OpenStudy (pinknabastak):

write an equation in vertex form of a parabola with zeros of -2 and 5 passing through (0,-15)

OpenStudy (pinknabastak):

please help!!!

OpenStudy (anonymous):

alright please note that the zeros are the x intercepts right ?

OpenStudy (pinknabastak):

yes

OpenStudy (anonymous):

so x-intercept form is just the factored form of standard form

OpenStudy (pinknabastak):

ok...

OpenStudy (anonymous):

alright so the zeros right do you remember how people get the zeros for an equation

OpenStudy (pinknabastak):

nope if i knew how to do the problem i would've done it myself

OpenStudy (anonymous):

alright I'm gonna have to teach you everything I know about solving polynomials with example problems get ready to get a math overload. REPEAT MATH OVERLOAD!! All people within a forty mail radius will be busted with math soon. You will thank me for this later. I not just gonna give the solution I goona teach it to you MUWHA!MUWHAHAHAHAHA!!

OpenStudy (pinknabastak):

hahaha ok...:) this sounds interesting..

OpenStudy (anonymous):

DO you know how to factor ?

OpenStudy (pinknabastak):

i used to its been two years since I've had to

OpenStudy (anonymous):

Now have to give you a problem that you have to try to factor atleast try okay .

OpenStudy (pinknabastak):

haha ok...im not the brightest lightbulb but ok

OpenStudy (anonymous):

Nah everyone is smart just had bad teachs all you need is a little moviation to do the problem and don't give ma negative attiude go solve this problem thinking your the smartest person the world.

OpenStudy (pinknabastak):

haha alright

OpenStudy (anonymous):

\[x^{2} + 10x+ 25\]

OpenStudy (anonymous):

Now factor and tell me how you did it

OpenStudy (pinknabastak):

its (x+5)(x+5)

OpenStudy (anonymous):

Do you know what the zeros of that problem is ?

OpenStudy (pinknabastak):

i did (x+ ?) (x+?) and then i just plugged in the numbers that fit... those were fairly easy

OpenStudy (pinknabastak):

-2 and 5

OpenStudy (anonymous):

(x+ 2)(x-5) - 15

OpenStudy (anonymous):

There's a helpful hint now put it in standard form.

OpenStudy (pinknabastak):

so thats in standard form?

OpenStudy (pinknabastak):

whys the 15 on the end there?

OpenStudy (anonymous):

Cause this is an equation it say it passing through - 15 that's the y intercept its always at the end the the problem the y intercept in when the x is at o and its just y

OpenStudy (anonymous):

but I need you to multiply (x+2)(x-5) for me to get standard form

OpenStudy (pinknabastak):

wait but i need vertex form..

OpenStudy (pinknabastak):

\[x^{2} - 3x-15\]

OpenStudy (anonymous):

add-15 to the equation

OpenStudy (pinknabastak):

so x^2 - 3x - 30

OpenStudy (anonymous):

 now use this formula to find the x

OpenStudy (displayerror):

pinknabastak, do you have Microsoft Excel on your computer?

OpenStudy (pinknabastak):

yes

OpenStudy (anonymous):

and yes once you find the x plug it in the the equation and get the why and I'll explain from there wait can we do this on microsoft excel tell me how

OpenStudy (displayerror):

You can easily check your solution on Microsoft Excel. I'll show you how, but first, I'll show you how to find the answer

OpenStudy (displayerror):

Recall that a parabola is of the general form \[ax^2 + bx + c = y\] We are given three points: two zeros and a point of intersection: (-2, 0), (5, 0), and (0, -15)

OpenStudy (pinknabastak):

its not nessicarly finding an answer but simply writing an equation

OpenStudy (pinknabastak):

yes thats all true

OpenStudy (displayerror):

That means we can plug those three points into the general formula to generate three separate equations to generate the equation of your parabola (from which you can then get the vertex form). Doing so, we get these three equations:

OpenStudy (displayerror):

For (-2, 0): \[4a - 2b + c = 0\] For (5, 0): \[24a + 5b + c = 0\] For (0, -15): \[c = -15\]

OpenStudy (displayerror):

You then have two equations and two unknowns--you can solve for a and b, then you will get your quadratic equation, from which you can then complete the square to arrive at the vertex form.

OpenStudy (displayerror):

You can also plug those three points into Microsoft Excel, plot the points (choose which data set is x, which data set is y). Then generate a polynomial trendline, display the equation, and you will get the plot here: http://i.imgur.com/vwNAFdK.png

OpenStudy (displayerror):

And using what rancd had mentioned before, you can find the intercept via the equation \[x = \frac{-b}{2a}\] and the corresponding y-value by plugging that x value into your final quadratic equation.

OpenStudy (pinknabastak):

so..... sorry i got kinda lost

OpenStudy (displayerror):

You are essentially finding the equation of your parabola by making use of the general quadratic equation \[f(x) = ax^2 + bx + c = y\] and the three given points

OpenStudy (pinknabastak):

huh??? so ax^2 +bx -15 = y

OpenStudy (displayerror):

Remember, you're given a coordinate, which contains both an x- and a y-value. Applying this idea to the coordinate (0, -15), then you have that x = 0 and y = -15. If you plug this into the general quadratic equation, you get \[a(0)^2 + b(0) + c = -15\] \[c = -15\] Therefore, \[f(x) = ax^2 + bx -15 = y\]

OpenStudy (pinknabastak):

so is that the answer?

OpenStudy (displayerror):

No, that is the first step in determining the values of your coefficients a, b, and c. Once you have those, you can complete the square and get the equation in vertex form.

OpenStudy (pinknabastak):

how do i complete the square?

OpenStudy (pinknabastak):

some one help! i just need the equation...

OpenStudy (ranga):

You have two roots: -2 and 5. That means (x-(-2)) and (x-5) are factors. So y = a(x+2)(x-5) where a is a constant that we need to find. (0,-15) is a point on the parabola. -15 = a(0+2)(0-5) -15 = -10a a = -15 / -10 = 3/2 So y = 3/2(x+2)(x-5) y = 3/2(x^2 -3x - 10) Put this in vertex form: y = 3/2 * { (x - 3/2)^2 - (3/2)^2 - 10 } = 3/2 * { (x-3/2)^2 -9/4 - 10 } y = 3/2 * { (x-3/2)^2 - 49/4 } y = 3/2(x-3)^2 - 147/8 is the equation of the parabola in vertex form.

OpenStudy (pinknabastak):

wow. Thank you zoo much @ranga its the only thing thats made sense to me in the past hour

OpenStudy (ranga):

Just let me double check for any mistakes in the calculations. Give me a sec.

OpenStudy (pinknabastak):

sorry one more... if you can.... write in factored form of parabola with vertex (5, 18/5) passing through (1, -14/5)

OpenStudy (pinknabastak):

ok

OpenStudy (ranga):

Typo in my last line. It should be: y = 3/2(x-3/2)^2 - 147/8

OpenStudy (pinknabastak):

haha thanks.. what about my other question?

OpenStudy (ranga):

The vertex form of a parabola is: y = a(x-h)^2 + k where (h,k) is the vertex. They have given you the vertex: (5, 18/5). So h = 5 and k = 18/5 y = a(x-5)^2 + 18/5 It passes through (1, -14/5). Put x = 1, y = -14/5 and find "a". -14/5 = a(1-5)^2 + 18/5 = 16a + 18/5 16a = -18/5 - 14/5 = -32/5 a= -32/(5*16) = -2/5 So y = -2/5(x-5)^2 + 18/5 = -2/5 * { (x-5)^2 - 9 } y = -2/5 * { x^2 -10x + 25 - 9 } = -2/5 * { x^2 - 10x + 16 } y = -2/5(x-8)(x-2) is the equation of the parabola in factored form.

OpenStudy (pinknabastak):

THANK YOU SOO MUCH!

OpenStudy (ranga):

You are welcome.